A chemist has on hand a supply of a 40% methyl alcohol solution and a 55% methyl alcohol solution. How much of each solution should be mixed to produce 2 L of a 43% solution?
I had tried singling out each variable but I'm not sure where I went wrong (my answers did not match up with the answer book, and I'd like to know how to get the correct answer)

Answer by Alan3354(69443)

A chemist has on hand a supply of a 40% methyl alcohol solution and a 55% methyl alcohol solution. How much of each solution should be mixed to produce 2 L of a 43% solution?
-------------------
x = 40%
y = 55%
---
x + y = 2 (total solution)
40x + 55y = 2*43 (total alcohol)